# BOOM: Benchmarking Out-Of-distribution Molecular Property Predictions of Machine Learning Models

### v1

@@ 2.3 Models
Graph Neural Networks
@@
GNNs are neural networks designed for learning on graph-structured data. Molecules and materials are represented as graphs of atoms and bonds, with 3D Euclidean space providing a natural molecular representation. As a result, message-passing neural networks (MPNNs) serve as the de facto backbone for deep learning-based molecular property prediction [19, 20]. Extensive work compares various GNN algorithms for this task. Instead of focusing on specific GNN variants, we examine the significance of architectural differences in our OOD task, emphasizing the relational inductive bias of molecular graphs and symmetries. (See Appendix 8.3).
@@
3D information and symmetries are fundamental to physical laws governing molecular behavior. Chemprop [21] serves as the baseline for a standard topological (2D) GNN. Additionally, we use three GNNs with topological, E(3) invariant, and E(3) equivariant learned models based on EGNN [22]. MACE is a popular E(3) equivariant GNN, which uses pair-wise distances for message passing and construction [23, 24]. Unlike EGNN, MACE also takes into account higher order interactions, potentially allowing for greater expressivity. To explore the effects of these symmetries, we test these five GNNs for our OOD tasks.
Hybrid Architectures
@@
Recently, we have seen an emergence of hybrid architectures that combine the inductive properties of GNNs and with the flexibility of the attention mechanism in Transformers. Graphormer [25] is a GNN-Transformer model that incorporates a graph-specific encoding mechanism to the input perform attention over structured data rather than sequences. Furthermore, Graphormer adds a bias term to the Query-Key product matrix to bias the attention to include bond information. We evaluate Graphormer-3D, a variant of Graphormer that incorporates inter-atomic distances to introduce 3D information to the attention mechanism. Finally, we also evaluate Equivariant Transformer (ET) [26], a 3D encoder-only transformer model that incorporate E(3) equivariance. Rather than inter-atomic distances, ET operates directly on 3D atomic coordinates.
2.4 Metrics
We consider a model to generalize well to the OOD test set if the RMSE prediction performance is comparable to that of the ID test set. However, most of the models tested were unable to reliably predict the exact numerical values of the OOD samples. Short of achieving strong OOD generalization, the next-best case is for the model to achieve a strong correlation on the OOD samples. Even if the model cannot accurately predict the correct numerical property values, achieving a strong correlation between the true and predicted property values is still useful for molecule discovery since the top-performing molecules will still be correctly identified. Altogether, we evaluate the OOD and ID RMSE prediction performance as a primary metric for measuring OOD generalization. We also evaluate the correlation on the OOD samples by calculating a binned $R^{2}$ value, which is the average value $R^{2}$ of the OOD samples in the lower and upper tails of the property distribution.
3 Related Work
Modern ML and DL models are designed for little discrepancy between training and test data distribution. OOD predictions present a key challenge for incorporating data-driven models into production pipelines where test time input may significantly shift from training data [27]. A possible solution to the OOD problem is to detect the distribution shift of input data away from the training distribution. OOD detection algorithms involve post-hoc or embedded updates to prediction models to detect when a sample is out of distribution. OOD detection has been approached through the lens of anomaly detection, uncertainty quantification, and open-set detection [6, 28, 29].
Rather than detecting distribution shifts, improving model robustness to unknown distribution shifts also addresses challenges posed by out-of-distribution data. Broadly speaking, the techniques for improving model robustness to distribution shifts can be divided into: (1) improved representation learning of features, (2) improved mapping of features to labels, and (3) improved optimization techniques [30].
In the domain of inorganic crystalline materials, Matbench [31] presents benchmarking on 13 tasks including the optical, thermal, electronic, thermodynamic, tensile, and elastic properties of inorganic materials. The Matbench tasks include 10 regression and 3 classification tasks based on the material composition and crystal structures. Follow-up studies utilized the Matbench datasets to evaluate the OOD performance of graph neural networks in both property and chemical structure space [32]. MatFold provides a convenient and systematic method to generate OOD test splits with respect to the structure of the material.[33] MatFold currently supports the creation of these OOD splits by structure, composition, chemical system, element, periodic group, periodic row, space group, point group, or crystal system. Our work differs from MatFold/Matbench in that i) we focus on OOD generalization in the property (y) space, instead of the input (x) space and ii) BOOM evaluates small molecule properties instead of inorganic crystalline materials.
In the small molecule space, the MoleculeNet is widely used for benchmarking molecular property prediction models. MoleculeNet consists of 17 small molecule prediction tasks, along with four splitting protocols: random, scaffold splitting, stratified splitting, and time splitting (test set consists of newest data). Segal et. al [34] also explore zero-shot extrapolation of molecular and material property prediction beyond the training data. Similar to our work, they also define OOD samples in the property space. They present bilinear transduction as a technique to enhance extrapolation capabilities. They evaluate on ESOL, Freesolv, Lipophilicity, and BACE binding from MoleculeNet for molecules. Bilinear transduction requires a descriptor-based model for anchor-based regression, while we focus on benchmarking existing deep-learning solutions.
4 Results
4.1 Survey of Model Architectures
Model Split HoF Density HOMO LUMO GAP ZPVE $R^{2}$ $\alpha$ $\mu$ $C_{v}$
Random ID 24.3453 0.0251 0.0061 0.007 0.0084 0.0012 52.0985 0.9441 0.6709 0.3785
Forest OOD 139.0728 0.1813 0.0304 0.0372 0.037 0.023 364.1557 8.4777 2.8977 3.3779
RT ID 28.1941 0.0224 0.0112 0.0135 0.0196 0.0034 95.8337 3.2265 1.2805 0.9291
OOD 496.6896 19403.8 54.6121 1.6136 0.0367 83.0197 374.4082 183239.6 2.9445 5.9453
ChemBERTa ID 22.86 0.0163 0.0068 0.0088 0.0103 0.0046 50.297 1.444 0.7134 0.4923
OOD 99.7253 0.1173 0.0245 0.0267 0.0315 0.0214 306.14 6.303 2.766 3.0175
MolFormer ID 9.3608 0.0175 0.0054 0.0054 0.0065 0.0015 40.9528 1.0544 0.6145 0.0826
OOD 93.6099 0.0924 0.0239 0.0248 0.0279 0.0213 283.0316 7.1359 2.2245 2.9328
Chemprop ID 15.68 0.0092 0.0041 0.0048 0.0058 0.0018 35.68 0.8305 0.55 0.3341
OOD 100.6 0.0551 0.0192 0.0187 0.0267 0.0129 234.73 4.7729 2.3 2.149
EGNN ID 9.3667 0.0076 0.0044 0.0044 0.0062 0.0013 20.2683 0.5655 0.4772 0.2504
OOD 17.7773 0.0296 0.0203 0.0223 0.029 0.0077 184.7985 5.4988 2.3559 2.1071
IGNN ID 19.7635 0.0086 0.0048 0.0051 0.0069 0.0015 30.5926 0.8374 0.5036 0.4018
OOD 24.5891 0.0261 0.0202 0.0195 0.0297 0.0075 184.1348 5.4229 2.4993 2.2091
TGNN ID 14.3601 0.0157 0.0054 0.0076 0.0372 0.0013 211.5326 0.7452 0.6371 0.3449
OOD 27.7216 0.0145 0.0178 0.0203 0.0129 0.0021 622.6762 2.8217 2.5002 0.5545
MACE ID 5.6737 0.0692 0.0121 0.0118 0.0172 0.0018 9.9994 0.3114 0.3731 0.1147
OOD 47.2462 0.1115 0.0279 0.0230 0.0392 0.0021 68.1583 1.5223 2.0991 0.1991
Graphormer ID 11.0218 0.008 0.0041 0.0043 0.0056 0.0002 33.0714 0.4326 0.6297 0.1932
(3D) OOD 30.6469 0.0255 0.0214 0.0226 0.0299 0.0141 264.556 4.4304 2.4939 1.4963
ET ID 50.4611 0.0079 0.0026 0.003 0.0042 0.0005 21.7465 0.3234 0.369 0.1296
OOD 108.1276 0.0341 0.0152 0.0137 0.0238 .0031 112.7228 3.0890 2.2832 0.9457
ModernBERT ID 12.8641 0.0113 0.0075 0.0095 0.0112 0.001 35.6684 1.1364 0.7342 0.5697
OOD 38.9671 0.0218 0.0226 0.0267 0.0351 0.0026 237.9504 2.6627 2.784 0.6373
Table 2: RMSE scores of all models on OOD and ID tasks. Best performing ID and OOD models are highlighted in Black and Blue respectively.The worst performing ID and OOD models are highlighted in Orange and Red respectively. The graph-based and hybrid models provide the best scores across nearly all tasks for OOD and ID splits. Numerical encoding issues greatly hamper RTs performance and result in large errors.
Figure 2: Normalized ID and OOD RMSE values for all models of interest for each property. RMSE values are normalized such that the best performing model has a value of 0 and the worst performing model has a value of 1. We omit the Regression Transformer model to remove the extreme values for better visualization. The models are sorted with respect to average RMSE, with the best model (lowest overall RMSE) appearing at the top of the heatmap and the worst model at the bottom.
First, we survey the performance of the most commonly used current machine learning models for predicting molecular properties (Table 2). The goal of this initial survey is to quantify the OOD performance that can be expected if one were to use these models in an ’out-of-the-box’ fashion without modifying the model architecture. However, we perform simple optimizations of the training parameters to achieve the highest possible accuracy for each task.
We visualize the relative performance of the models on ID and OOD tasks in Figure 2, wheresas Table 2 provides the complete results with all models. We do not find any model that clearly outperforms the others on either ID or OOD performance across all tasks. Among all tested models, the Equivariant Transformer (ET) model achieves the best overall ID performance - achieving the lowest ID RMSE on 4 out of 10 tasks. For OOD prediction, MACE achieves top performance on 5 out of 10 tasks, and ET achieves top performance on 3 out of 10 tasks. We similarly find that ET achieves the best overall ID performance and MACE achieves the best OOD performance when evaluated according to the binned $R^{2}$ metric (see Table 4). We note that the Regression Transformer model is unique among the tested models in that property prediction is performed via autoregressive numerical token prediction. The large OOD RMSEs noted by Regression Transformer were found to arise from inaccuracies in this autoregressive numerical token generation, for example, predicting ’00913’, for a true value of ’0.913’. Figure 3 shows a common mode of failure for OOD predictions for most models (see parity plots for all models tested in Appendix 9). We find that models performing poorly on OOD splits overwhelmingly produce an S-shaped parity plot. The models are therefore capable of clustering OOD samples together but are unable to extend the prediction region beyond the training data. Such S-shaped behavior is a known failure case that arises when models learn short-cut features that maximize ID performance, but fail to generalize to OOD data [35].
Furthermore, we notice MoLFormer, one of the largest models in our test suite, achieves near state of the art performance on HoF and $C_{v}$, greatly outperforming other models, including similar Transformer models like ChemBERTa and RT. But ChemBERTa and MolFormer achieve similar results on OOD for both tasks. We notice a similar trend where ID performance is not necessarily correlated with OOD performance. Considering the size of our datasets, we believe large models may be able to overfit to the ID space, while achieving subpar generalization. This suggests the common strategy of pre-training on large datasets and fine-tuning on niche domains may have pitfalls for OOD samples.
@@ 4.1 Survey of Model Architectures
Figure 3: Representative parity plots illustrating the OOD (blue) and ID (orange) test predictions on the 10K solid heat of formation dataset. (Left) MoLFormer predictions on this task exhibit poor OOD performance with weak correlation on the OOD samples. (Right) MoLFormer displays strong OOD performance on the QM9 zero-point vibrational energy task.
We also provide a task-centric view of our results in Fig. 4. We highlight the mean for each split with a horizontal bar in Fig. 4. The larger the difference between the ID and OOD bars, the higher the discrepancy between ID and OOD performance. As expected, ID performance is better than OOD performance for all model-task pairs. We highlight that good OOD performance is achievable in certain tasks, such as HoF, Density, ZPVE, and $C_{v}$. Some models achieve near-ID-level performance on these particular tasks. However, Fig. 4 also highlights particular tasks (HOMO, LUMO, Gap, and Dipole Moment($\mu$)) where no models achieve good OOD performance. Since all these properties are related to the electronic structure of molecules, we hypothesize that the inability of any model to generalize well in these tasks is due to the lack of explicit electronic structure information in their molecular representations. It is also important to note that for properties such as $C_{v}$, although most models achieve similar ID $R^{2}$ values, there is a large variance in OOD binned $R^{2}$ values- further highlighting the importance of performing OOD performance evaluations.
Figure 4: We present binned $R^{2}$ scores for OOD and standard $R^{2}$ scores for ID on each task for all models. The orange and blue bars indicate the performance averaged across all models for ID and OOD, respectively. Nearly all models have significant discrepancies between ID and OOD performance, but some models can reach ID-level accuracy. We observe that OOD performance is highly task-dependent.
4.2 Impact of Pre-Training
In this section, we seek to understand to what extent the large scale pretraining of existing chemical foundation models improves their OOD generalization performance. In analogy to the work in language models, these chemical foundation models are trained on vast pretraining datasets of billions of molecules with the hope that this pretraining will provide a general understanding of fundamental chemical principles. We benchmark ChemBERTa and MoLFormer (both MLM pretraining) and Regression Transfomer (PLM pretraining) to understand how the choice of pretraining tasks impacts OOD performance. Notably, the original reports of all three of these foundation models showed that that this large-scale language pretraining strategy can achieve SOTA performance on in-distribution molecular property prediction tasks[11, 13], but did not perform evaluations of the OOD performance.
Across all three foundation models, we see a consistent improvement of ID performance across the majority of tasks Figure 11 . Averaged across all 10 tasks, the pretrained models show a sizable improvement in ID RMSE due to pretraining (31% for ChemBERTa, 35% for MoLFormer and 12% for Regression Transformer). We note that these results are consistent with the findings in their original reports. For example, the MoLFormer paper found a 29% reduction in MAE ID performance on the QM9 dataset due to MLM pretraining, whereas Regression Transformer reported up to a 52% reduction in predicting ID QED RMSE from optimizing the pretraining objective. A similar ablation study was not performed in the original ChemBERTa paper.
To explore if strong OOD generalization can be achieved through alternative pretraining tasks, we also explore pretraining on a supervised property prediction task. First, we perform supervised pretraining of a Chemprop model on the entirety of one of the eight QM9 property datasets. This pretrained model is then finetuned on only the training set of one of the other seven QM9 property dataset. This isolates only the property to be OOD, as the model has seen all the molecules in another context. We can see in Fig. 6 OOD performance is only improved when the pretraining task is sufficiently related to the downstream finetuning task. Specifically, across all eight datasets, we see a significant reduction in the OOD RMSE (compared against the performance without pretraining) when the Pearson correlation coefficient between the pretraining and finetuning tasks is less than 0.35. This result may explain why the masked language modeling pretraining used in current chemical foundation models resulted in worse OOD performance (Fig. 5).
4.3 Hyperparameter Optimization
Figure 7: OOD Performance of the Chemprop MPNN model when using default hyperparameters and the best performing OOD hyperparameters. The best OOD hyperparameters are determined according to the maximum OOD test RMSE for each property. For each property, we randomly sampled 50 values of the following hyperparameters: the number of message-passing steps; the number of layers in the feed-forward network; the dropout probability; the hidden dimension in the message-passing step; and the hidden dimension in the feed-forward network.
As shown in Fig. 7, we compare the OOD performance of Chemprop when using the default hyperparameters and hyperparameters that have been optimized to maximize OOD performance. Overall, we do not find that hyperparameter optimization can provide meaningful improvements to OOD performance. While we see a noticeable reduction in the OOD RMSE in relatively simple properties such as density(-60%), heat of formation(-23%) and ZPVE(-50%) following hyperparameter tuning, the models are still unable to generalize significantly beyond the training regime. It may not solve the problem, but for certain properties, hyperparameter optimization with respect to OOD improved over the default model by 60%, without any significant decrease to the ID performance. The results here highlight that OOD performance should be considered as an important evaluation criterion for future model optimization to ensure that models strike a balance between ID performance and OOD generalization.
4.4 Representation

### v2

@@ B.1 Model Summary
Multilayer Perceptron Multilayer Perceptron RDKit Molecular Descriptors N/A 153k
@@
GotenNet GNN {Atom, Bond, Pair-wise Distances} E(3)-equivariant + permutation 6M
@@
GeoFormer Hybrid {Atom, Bond, Pair-wise Distances } E(3)-invariant + permutation 47.1M
@@
Model Device Runtimes (10k) [seconds/epoch] Runtimes (QM9) [seconds/epoch]
Random Forest 2.4 Ghz 8-core Intel Core i9 0.5 6
RT V100 632 10275
ChemBERTa L40 13 165
MoLFormer H100 14 72
Chemprop H100 10 23
EGNN L40 10 115
IGNN L40 10 115
TGNN L40 10 115
MACE H100 70 165
Graphormer AMD MI300A 6 19
ET L40 10 75
Geoformer H100 60 230
GotenNet A100 20 120
ModernBERT L40 15 165
MLP H100 0.009 0.5
Table 4: Runtimes for all models used in BOOM on the 10k and QM9 datasets.
B.2 RDKit Featurizer
The RDKit Featurizer, as implemented in the Deepchem package,[Ramsundar-et-al-2019] consists of 125 chemically-informed features (such as molecular weight and number of valence electrons), as well as 86 features describing the fraction of atoms that belong to notable functional groups such as alcohols or amines.
B.3 Transformers
Transformers, including large language models (LLMs), have revolutionized language modeling and vision tasks and have gained popularity in scientific regimes. We choose four representative models to cover the major archetypes of transformer models: MoLFormer  [MolFormer2022], ChemBERTa  [chithrananda2020chemberta], Regression Transformer  [born2023regression], and ModernBERT  [modernBERT].
We choose three representative models to cover the major archetypes of transformer models. MoLFormer  [MolFormer2022] is an encoder-decoder model with a T5  [t5] backbone originally trained on PubChem. ChemBERTa  chithrananda2020chemberta is an encoder-only model with a BERT  [bert] backbone trained on PubChem. Finally, we also use Regression Transformer  [born2023regression], an XLNet-based  [yang2019xlnet] model that is capable of both masked language modeling as well as autoregressive generation.
B.3.1 ModernBERT
We also evaluate ModernBERT, a state-of-the-art (SOTA) encoder-only model with architectural improvements such as rotary positional embeddings [rope], pre-normalization, and GeGLU activation layers [glu]. Along with different architectures, we also investigate the effects of different pre-training and tokenization schemes in our experiments. The training details are presented in Appendix  C.2.
B.4 GNNs
@@
GNNs are neural networks designed for learning on graph-structured data. Molecules and materials are represented as graphs of atoms and bonds, with 3D Euclidean space providing a natural molecular representation. As a result, message-passing neural networks (MPNNs) serve as the de facto backbone for deep learning-based molecular property prediction [schnet, orbnet]. Extensive work compares various GNN algorithms for this task. Instead of focusing on specific GNN variants, we examine the significance of architectural differences in our OOD task, emphasizing the relational inductive bias of molecular graphs and symmetries.
@@
3D information and symmetries are fundamental to physical laws governing molecular behavior. Chemprop [heid2023chemprop] serves as the baseline for a standard topological (2D) GNN. Additionally, we use three GNNs with topological, E(3) invariant, and E(3) equivariant learned models based on EGNN [EGNN]. MACE is a popular E(3) equivariant GNN, which uses pair-wise distances for message passing and construction [batatia_mace_2023, kovacs_evaluation_2023]. Unlike EGNN, MACE also takes into account higher-order interactions, potentially allowing for greater expressivity. To explore the effects of these symmetries, we test these five GNNs for our OOD tasks.
Symmetries are inherent to the physical laws that dictate molecular properties. Algebraically, they are represented as groups, where each element corresponds to a transformation. For non-chiral molecules, the E(3) group, encompassing rotations, translations, and reflections, is key. Chiral molecules require the SE(3) subgroup, which excludes reflection. Since molecular properties remain invariant under these transformations, learned structure-to-property functions should obey the same symmetries.
GNNs naturally encode these symmetries. MPNNs enforce permutation-invariant message aggregation, making models permutation-invariant. Geometric deep learning models can extend this by enabling molecular representations in 3D space, ensuring networks are invariant or equivariant to geometric transformations. Invariance implies properties remain unchanged after transformation, while equivariance means vector properties transform consistently with applied transformations. Here, we provide rigorous definitions.
For completeness, we reproduce the GNN formulation from [EGNN]. For a given GNN with node features $h^{(l)}_{i}$ are the features of the $i$-th node for $l$-th layer. $b_{ij}$ are the edge-features between two connected nodes $i$ and $j$ such that $j\in\mathcal{N}_{i}$. The neighborhood $\mathcal{N}_{i}$ is the set of nodes connected to node $i$. $W^{(l)}$ is a learnable projection matrix of layer $l$.
Topological GNN:
$h^{(l+1)}_{i}=h^{(l)}_{i}W^{(l)}+\sum_{j\in\mathcal{N}_{i}}\theta(b_{ij},h^{(l)}_{i},h^{(l)}_{j})$ (3)
Where $\theta(\cdot)$ is a learnable function of the bond and node features, shared between all node pairs.
Invariant GNN:
$h^{(l+1)}_{i}=h^{(l)}_{i}W^{(l)}+\sum_{j\in\mathcal{N}_{i}}\theta(b_{ij},h^{(l)}_{i},h^{(l)}_{j})+\sum_{j\neq i}\phi(r_{ij},h^{(l)}_{i},h^{(l)}_{j})$ (4)
[…]
Equivariant GNN:
$h^{(l+1)}_{i}$ $=h^{(l)}_{i}W^{(l)}+\sum_{j\in\mathcal{N}_{i}}\theta(b_{ij},h^{(l)}_{i},h^{(l)}_{j})+\sum_{j\neq i}\phi(r^{(l)}_{ij},h^{(l)}_{i},h^{(l)}_{j})$ (5)
$x^{(l+1)}$ $=x^{(l)}+\sum_{j\neq i}\left(\frac{x^{(l)}_{i}-x^{(l)}_{i}}{r^{(l)}_{ij}+\xi}\right)\psi(r^{(l)}_{ij},h^{(l)}_{i},h^{(l)}_{j})$ (6)
[…]
As we can see Eq. 5 is equivalent to Eq. 4 but with a per-layer coordinate update. Furthermore, Eq. 4 is equivalent to Eq. 3 but with an additional term dependent on the pairwise distances $r_{ij}$.
B.4.1 Readout Function
The readout function, $\mathcal{R}$ of a GNN aggregates the node-level information on the graph and combines them to get a graph-level output. The readout function can be any permutation invariant function such that, $\mathcal{R}:\mathbb{R}^{|\mathcal{V}\times F|}\rightarrow\mathbb{R}^{K}$, where $F$ is the per-vertex feature dimension, and $K$ is the output dimension ($K=1$ in the case of regression). The flexibility in the readout function can be used to provide target-specific inductive bias such as using a summing over the vertices for extensive properties while taking the mean output for the vertices for intensive properties.
MACE and ET use modified readout functions for some properties, such as $\mu$, while we are using the unmodified readout function. We have not had success modifying the readout function as described in their publication, but we are working with the authors to replicate their results. We plan on investigating this further.
B.5 Hybrid Architectures
@@
Recently, we have seen an emergence of hybrid architectures that combine the inductive properties of GNNs and with the flexibility of the attention mechanism in Transformers. Graphormer [graphormer] is a GNN-Transformer model that incorporates a graph-specific encoding mechanism to the input perform attention over structured data rather than sequences. Furthermore, Graphormer adds a bias term to the Query-Key product matrix to bias the attention to include bond information. We evaluate Graphormer-3D, a variant of Graphormer that incorporates inter-atomic distances to introduce 3D information to the attention mechanism. Finally, we also evaluate Equivariant Transformer (ET) [torchmd], a 3D encoder-only transformer model that incorporates E(3) equivariance. Rather than inter-atomic distances, ET operates directly on 3D atomic coordinates.
Appendix C Training Details
C.1 General Training
Across all models, we generally hold out $10\%$ of the training data for hyperparameter selection. As we had multiple different types of models, we started with publicly available settings for the starting hyperparameters (as noted below), but also performed hyperparameter sweeps (with grid search) on the non-architectural components, such as learning rate and training steps. We do not update architectural details to match the use case of practitioners using off-the-shelf models. The GNN ablation uses the architecture detailed here,EGNN and training instructions listed in the Appendix of that work. The baseline models (Random Forest and MLP), use model hyperparameters previously reported.yang2019analyzing, moleculenet
C.2 Transformer Fine-tuning Details
ChemBERTa and MoLFormer models are pre-trained with a masked language modeling (MLM) task and the Regression Transformer is pretrained with a permutation language modeling (PLM) task. During MLM pretraining, a predetermined fraction of the SMILES string of the molecule is masked and then predicted by the model. The Regression Transformer foundation model uses a PLM pretraining task, which seeks to autoregressively predict masked tokens from a permuted sequence of both SMILES and property tokens.
For Regression Transformer and ChemBERTa, the models without pretraining are initialized with random weights, whereas the MoLFormer model without pretraining is loaded directly from the provided checkpoint saved at the beginning (0th iteration) of pretraining. For all three models, the pretrained models are initialized from the provided model checkpoints, before finetuning on each of the 10 downstream OOD tasks. Both the pretrained and scratch models are fine-tuned according to the same learning schedule hyperparameters.
C.3 Chemprop Pretraining Details
For the experiments highlighted in Figure 5, we first train all model weights of the Chemprop model (v1.4.0) for 30 epochs on the entirety of one of the eight QM9 property datasets (133,885 training examples). Then, this model is finetuned for an additional 30 epochs on only the train split (see 2) of one of the other seven QM9 property datasets. During this finetuning step, the model parameters of the message-passing neural network portion of Chemprop are frozen. All other model hyperparameters are the Chemprop defaults and are provided in the Github repo.
C.4 Chemprop Hyperparameter Optimization
To understand to what extent hyperparameter optimization can affect OOD performance, we first train the Chemprop model (v1.4.0) with all default hyperparameters on all 10 BOOM datasets. Then, we train Chemprop on each of the 10 BOOM datasets with 50 independent, random choices of model hyperparameters (i.e. with 50 random seeds). The tuned model hyperparameters are the message passing depth, sampled from between 2-6 layers, the fraction of dropout in the neural network sampled between 0-0.40 with an increment of 0.05, the number of feed-forward layers sampled from between 1-3 layers, and the size of the hidden layers, sampled between 300-2400 with an increment of 100.
C.5 Additional Plots
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Chemprop MPNN Performance with Data Augmentation and QM9 OOD tasks
Figure 10: OOD Performance of chemical foundation models (ChemBERTA, MoLFormer and Regression Transformer) with and without pretraining.The performance of Regression Transformer on the QM9 dipole moment and HOMO-LUMO Gap properties are omitted due to the inability of the scratch Regression Transformer model to converge on these properties.
C.6 Tabulated results